# Monitoring Latent World States in Language Models with Propositional Probes

### v1

@@ 5 Propositional Probes
In this section we construct and evaluate propositional probes. To construct propositional probes, we first train 4 domain probes (Sec. 5.1), then compose them together using the binding similarity metric (1) obtained from the binding subspace (Sec. 5.2). We then evaluate our probes on the three datasets synth, para, and trans. Despite being only trained on synth, propositional probes perform comparably with a prompting skyline on all three datasets (Sec. 5.3).
@@ 5.1 Domain probes
For every domain $\mathcal{D}_{k}$, we train a probe $P_{k}$ that classifies the activation at each position $Z_{s}$ into either a value $P_{k}(Z_{s})\in\mathcal{D}_{k}$ in the domain or the null class $\bot$. We use the activation at a particular layer $l$ as the input of the probe. We describe later how $l$ is chosen. The probe then has the type $P_{k}:\mathbb{R}^{d_{\text{model}}}\rightarrow\mathcal{D}_{k}\cup\{\bot\}$.
We parameterize the probe with $|\mathcal{D}_{k}|$ vectors, $u_{k}^{(0)},\dots u_{k}^{(|\mathcal{D}_{k}|-1)}\in\mathbb{R}^{d_{\text{model}}}$ as well as a threshold $h_{k}\in\mathbb{R}$. Each vector can be thought of as a direction in activation space corresponding to a value in the domain. The classification of the probe is simply the value whose vector has the highest dot product with the activation. When all vectors have a dot product smaller than the threshold, the probe classifies $\bot$ instead. Formally,
@@
To learn the vectors, we generate a dataset of activations and their corresponding values. Then, we set each vector to the mean of the activations with that input. Then, we subtract each vector with the average vector $\frac{1}{|\mathcal{D}_{k}|}\sum_{i}u_{k}^{(i)}$. This can be seen as a multi-class generalization of the so-called difference-in-means probes  [24].
To generate this dataset, we use synthetic templates. However, this only provides context-level supervision: we know that the activations in the context collectively represent a certain value in the domain, but we do not know which activations represent the value and which represent $\bot$. Thus, we have to assign the activations at each token position with a ground-truth label.
To do so, we use a Grad-CAM-style attribution technique  [38] similar to that used by Olah et al. [39]. Broadly speaking, we backpropagate through the model to estimate how much the activation at each layer/token position contributes towards the model’s knowledge of the content. This identifies both the token position that carries value information and the layers that are the most informative.
The attribution results indicate that the middle layers at last token position are the most informative. We thus choose layer $l=20$ (out of 40 layers). In Appendix G we discuss the attribution results further. In Sec. 5.3 we evaluate the accuracy of domain probes on the synth dataset on which they are trained, as well as the para and trans datasets.
@@ 5.2 Propositional probes
With the binding similarity (Eq. 1) obtained from the binding subspace, we can compose domain probes with a simple lookup algorithm. We first identify all the names mentioned in the context with the name domain probe $P_{0}$. Then, for every other domain probe $P_{k}$, we identify the values it picks up in the context, and for each of these values we look for the best matching name to compose together. The pseudocode is described in Alg. 1.
@@
1:Domain probes $\{P_{k}\}_{k}$ and binding form $H$.
@@
5.3 Evaluations of domain and propositional probes
Figure 6: Left: Accuracy of domain probes on synth. Middle: Exact match accuracy and Jaccard index of probing and prompting. “DAS-1”, “DAS-50”, and “random” are ablations of propositional probes where we use the 1-dim and 50-dim DAS subspace or a random 50-dim subspace as the binding subspace. Right: Exact match accuracy and Jaccard index of probing and prompting on prompt-injected (p) and backdoored (ft) settings. All: Error bars indicate standard error.
In this section we evaluate propositional probes on the three datasets, synth, para, and trans. The constituents of propositional probes, domain probes and the binding subspace, are obtained from templated data, and we thus view para and trans as evaluating out-of-domain generalization. We first evaluate each of the four domain probes in isolation, and then evaluate the propositional probe.
Domain probes   We evaluate the exact-match accuracy of each of the four domain probes at predicting the correct set of values that are in the context for the three datasets (Fig. 6 left). Each context contains $N=2$ values, and domain probes have to get both right. Despite being trained only on a synthetic dataset, we find that domain probes generalize to the paraphrased dataset para, and even the Spanish dataset trans.
@@ 5.3 Evaluations of domain and propositional probes
Propositional probes   We evaluate the Jaccard index and exact-match accuracy between the ground-truth propositions and the set of propositions returned by the propositional probe (or prompting skyline). Specifically, for an input context, let the set of propositions returned by the propositional probe (or prompting) be $A$, and the ground-truth set of propositions be $B$. The exact-match accuracy is the fraction of contexts for which $A=B$, and the Jaccard index is the average value of $|A\cap B|/|A\cup B|$. Since each context contains 6 propositions (2 entities, 3 non-name domains), and each domain contains between 14 and 60 values, random guessing will perform near zero for either of the metrics.
@@
We compare propositional probes against a prompting skyline that iteratively asks the model questions about the context. First, we ask for the number of names. Then, we ask for the names. For each name, we ask for the associated value for every non-name domain (e.g. “What is the occupation of John?”), and select the value in the domain with the highest log probability. High performance at the prompting skyline validates the assumption that the model understands the context well enough to answer questions about it. In addition, we evaluate ablations of our propositional probes in which the Hessian binding metric is replaced with the dot-product following a projection into the 50-dimensional and 1-dimensional DAS subspaces (“DAS-50” and “DAS-1”), and a 50-dimensional random subspace (“random”) respectively.
@@
We find that propositional probes do comparably with the prompting skyline in terms of Jaccard index for all three datasets (Fig. 6 middle), but do significantly worse for terms of exact-match accuracy for trans. We also note that the accuracies of the domain probes are strict upper bounds on the accuracy of the propositional probe, which suggests a lot of the drop in accuracy in trans could be due to the food probe generalizing less well to the Spanish trans dataset.
6 Applications

### v2

@@ 4 Propositional probes
In this section we describe the overall architecture of propositional probes. At a high level, propositional probes are composed of domain probes, one for each domain, and the outputs of the domain probes are composed to form propositions using the binding similarity metric. We construct domain probes in Sec. 4.1, compose them to form propositional probes in Sec. 4.2, but defer the construction of the binding similarity metric to Sec. 5. We make code available at https://github.com/jiahai-feng/prop-probes-iclr
@@ 4.1 Domain probes
For every domain, we train a domain probe that linearly classifies activations at individual token positions into either a value in the domain, or a null value $\bot$, indicating that none of the values is represented. For example, Fig. 1 (left) shows the outputs of the name probe and the country probe. Formally, for every domain $\mathcal{D}_{k}$, we train a probe $P_{k}:\mathbb{R}^{d_{\text{model}}}\rightarrow\mathcal{D}_{k}\cup\{\bot\}$. We use the activation at a particular layer $l$ as the input of the probe. We describe later how $l$ is chosen.
We parameterize the probe with $|\mathcal{D}_{k}|$ vectors, $u_{k}^{(0)},\dots u_{k}^{(|\mathcal{D}_{k}|-1)}\in\mathbb{R}^{d_{\text{model}}}$ and a threshold $h_{k}\in\mathbb{R}$. Each vector is a direction in activation space corresponding to a value in the domain. The classification of the probe is simply the value whose vector has the highest dot product with the activation, or the null value $\bot$ instead if all the dot products are smaller than the threshold. Formally,
@@
To learn the vectors, we generate a dataset of activations and their corresponding values. Then, we set each vector to the mean of the activations with that input. Then, we subtract each vector with the average vector $\frac{1}{|\mathcal{D}_{k}|}\sum_{i}u_{k}^{(i)}$. This can be seen as a multi-class generalization of the so-called difference-in-means probes  (Mallen & Belrose, 2023).
We collect this dataset of activations from the synth dataset. However, this only provides context-level supervision: we know that the activations in the context collectively represent certain values in the domain, but we do not know which activations represent the value and which represent $\bot$. Thus, we have to assign the activations at each token position with a ground-truth label.
To do so, we use a Grad-CAM-style attribution technique  (Selvaraju et al., 2017) similar to that used by Olah et al. (2018). Broadly speaking, we backpropagate through the model to estimate how much the activation at each layer/token position contributes towards the model’s knowledge of the lexical information, which estimates the saliency of both the layer and token position.
The attribution results indicate that the middle layers at last token position are the most informative. We thus choose layer $l=20$ (out of 40 layers). We discuss the attribution further in Appendix G.
@@ 4.2 Propositional probes
Our propositional probes compose the outputs of constituent probes with the binding similarity metric using a simple lookup algorithm. We first describe the binding similarity metric, and then describe the algorithm for propositional probes.
We expect the binding similarity metric to be greater for activations that are bound together than activations that are not. In the example in Fig. 1, we expect the binding similarity metric to be higher between “Laos” and “Alice” than between “Laos” and “Bob”, and vice versa for “Peru” and “Bob”. Concretely, the binding similarity metric $d(Z_{s},Z_{t})$ is a real-valued function that describes how strongly bound the activations $Z_{s}$ and $Z_{t}$ are to each other. In Sec. 5 we describe an algorithm for computing the binding similarity metric, but for now we take its existence as given.
Our propositional probes use the binding similarity metric to decide how the predicted attributes should be bound to the predicted entities. Specifically, we first identify all the names mentioned in the context with the name domain probe $P_{0}$. Then, for every other domain probe $P_{k}$, we identify the values it picks up in the context, and for each of these values we select the name with the highest binding similarity metric to compose together. The pseudocode is described in Alg. 1.
@@
1:Domain probes $\{P_{k}\}_{k}$ and binding similarity metric $d(\cdot,\cdot)$.
@@
5 Binding subspace
In this section, we present a Hessian-based algorithm for identifying the binding subspace, from which we construct the binding similarity metric used in the propositional probes. We first review background information on the binding subspace (Sec. 5.1), before describing our algorithm for identifying the binding subspace and constructing the binding similarity metric (Sec. 5.2). Lastly, we show that the resultant binding subspace is causally mediating (Sec. 5.3).
5.1 Binding subspace background
The binding subspace arises from a decomposition of internal activations into lexical and binding vectors (Feng & Steinhardt, 2023). While researchers have long noted that language model activations often comprise of lexical vectors that linearly encode lexical concepts, Feng & Steinhardt (2023) recently observed that activations can also contain binding vectors that bind lexical concepts in one activation with lexical concepts in another.
Specifically, suppose the input describes two entities $E_{0},E_{1}$ (e.g. “Alice”, “Bob”) with corresponding attributes $A_{0},A_{1}$ (e.g. “Laos”, “Peru”) (see Fig. 1). Let $Z_{E_{k}}$ and $Z_{A_{k}}$ for $k=0,1$ be the activations for $E_{k}$ and $A_{k}$ respectively. Then, Feng & Steinhardt (2023) observed that the activations can be decomposed into lexical vector representations $f_{E}(E_{k})$ and $f_{A}(A_{k})$ of entities and attributes, and binding vectors $b_{E}(k)$ and $b_{A}(k)$,
$Z_{E_{k}}=f_{E}(E_{k})+b_{E}(k),\quad Z_{A_{k}}=f_{A}(A_{k})+b_{A}(k),$
so that interventions that modify $Z_{E_{k}}\leftarrow Z_{E_{k}}+b_{E}(k^{\prime})-b_{E}(k)$ for $k^{\prime}\neq k$, will switch the bound pairs (e.g. from $\{(E_{0},A_{0}),(E_{1},A_{1})\}$ to $\{(E_{0},A_{1}),(E_{1},A_{0})\}$), and likewise for $Z_{A_{k}}$.
However, the techniques in prior work could not identify the binding subspace (i.e. the subspace binding vectors live in), or the geometric relationship between the binding vectors $[b_{E}(k),b_{A}(k)]$. In particular, they relied on estimating differences in binding vectors $\smash{\Delta_{E}\triangleq b_{E}(1)-b_{E}(0)}$, $\smash{\Delta_{A}\triangleq b_{A}(1)-b_{A}(0)}$. To do so they computed the difference $Z_{E_{1}}-Z_{E_{0}}$ across 200 contexts with different values of $E_{0}$ and $E_{1}$, with the hope that the content dependent vectors $f_{E_{0}},f_{E_{1}}$ cancelled out (likewise for $\Delta_{A}$). This enabled the bound pair switching intervention described in the preceding paragraph, but did not reveal how the binding vectors are geometrically related.
In this work, we hypothesize that there is low rank linear relationship between binding vectors that correspond to each other, i.e. there exists a low-rank binding matrix ${\bm{H}}$ under which related entity/attribute binding vectors have high norms, so that $b_{E}(k)^{\top}{\bm{H}}b_{A}(k^{\prime})$ is high only when $k=k^{\prime}$. We further posit that ${\bm{H}}$ is orthogonal to the subspaces spanned by the lexical vectors $f_{E}(E_{k}),f_{A}(A_{k})$, so that $Z_{E_{k}}^{\top}{\bm{H}}Z_{A_{k^{\prime}}}\approx b_{E}(k)^{\top}{\bm{H}}b_{A}(k^{\prime})$. In Sec. 5.2 we describe a Hessian-based algorithm to identify ${\bm{H}}$, and construct a symmetric metric $d(Z_{k},Z_{l})$ that describes whether $Z_{k}$ is bound to $Z_{l}$, which is used ultimately by propositional probes to compose domain values together.
5.2 Hessian-based algorithm
To motivate the Hessian-based algorithm, we briefly discuss a straightforward approach to estimate the binding matrix ${\bm{H}}$ and a key shortcoming. A simple way to estimate ${\bm{H}}$ is to collect a dataset of model activations, together with ground-truth labels of which activations are bound together. Then, one can obtain ${\bm{H}}$ as either the result of bilinear regression, or with a more sophisticated loss involving causal interventions (Geiger et al., 2024) [^1] We implement Distributed Alignment Search as a baseline in Sec. 5.3. However, in either case, the estimated ${\bm{H}}$ is limited because it can only ever capture variations on binding vectors present in the initial dataset; in our setting where we train on a simple dataset and test on hard, diverse datasets, we need a method that captures structure inherently present in the model, and not just in the dataset.
At a high level, the Hessian-based approach approximates the model with its Hessian, which characterizes not just the model’s behavior on binding vectors in the present input, but in all directions in the activation space. Specifically, suppose we have two activations $Z_{x},Z_{y}$ corresponding to an entity and an attribute which are initially unbound, and we perturb the first in direction $x$ and the latter in direction $y$. Suppose further we have a function $F(x,y)$ that measures the binding strength between $Z_{x}+x$ and $Z_{y}+y$. The binding strength should increase only if directions $x$ and $y$ align under ${\bm{H}}$. In fact, if $F(x,y)$ is bilinear, we can recover ${\bm{H}}\approx\nabla_{x}\nabla_{y}F(x,y)$, up to a scaling coefficient.
Figure 3: Overview of Hessian-based algorithm. 1) Activations for “Alice” and “Laos” are bound because their binding vectors (horizontal) align under binding matrix ${\bm{H}}$, likewise for “Bob” and “Peru”. 2) Ablate binding information by setting binding vectors to midpoints. 3) Perturb activations with ±$\pm x$ and ±$\pm y$; binding is recovered in figure because $x$ and $y$ are aligned. 4a, 4b) To compute binding strength $F(x,y)$, append query strings and measure the probability of correct next token.
To instantiate this, we need a way of creating unbound activations, and also of measuring binding strength. To create unbound activations, we construct activations that are initially bound and subtract away the binding information so that the binding vectors become indistinguishable. Specifically, consider the context with two propositions $\{(E_{0},A_{0}),(E_{1},A_{1})\}$ discussed in Sec. 3. To ablate binding information in $Z_{E_{k}}$ and $Z_{A_{k}}$, we add $0.5\Delta_{E}$ and $0.5\Delta_{A}$ to $Z_{E_{0}}$ and $Z_{A_{0}}$, and subtract the same quantities from $Z_{E_{1}}$ and $Z_{A_{1}}$. This moves the binding vectors to their midpoint and so should cause them to be indistinguishable from each other (Step 2 in Fig. 3).
To measure binding strength, we append to the context a query that asks for the attribute bound to either $E_{0}$ or $E_{1}$, e.g. “Which country does $E_{0}$/$E_{1}$ live in?”. We measure the probability assigned to the correct answer ($A_{0}$ and $A_{1}$ respectively), and take the average over the two queries as the measure of binding strength $F(x,y)$, where $x$ and $y$ are the perturbations added to the unbound entity and attribute activations. Therefore, when all activations are unbound ($x=y=0$), the model takes a random guess between the two attributes, and so $F(0,0)=0.5$. If $x$ and $y$ are aligned under $H$, we expect $F(x,y)>0.5$ (Step 3 in Fig. 3).
The matrix ${\bm{H}}$ is obtained as the second-derivative $\nabla_{x}\nabla_{y}F(x,y)$. For tractability, we parameterize $x$ and $y$ so that they are shared across layers, i.e. $\smash{x,y\in\mathbb{R}^{d_{\text{model}}}}$ instead of $\smash{\mathbb{R}^{d_{\text{model}}\times n_{\text{layers}}}}$. Further, to obtain the low-rank binding subspace from ${\bm{H}}$, we take its singular value decomposition ${\bm{H}}={\bm{U}}{\bm{S}}{\bm{V}}^{\top}$. We expect the binding subspace to be the top $k$-dimensional subspaces of ${\bm{U}}$ and ${\bm{V}}$, ${\bm{U}}_{(k)},{\bm{V}}_{(k)}\in\mathbb{R}^{d_{\text{model}}\times k}$, for a relatively small value of $k$.
Binding similarity metric   To turn the binding subspace ${\bm{U}}_{(k)}$ into a symmetric similarity metric between two activations $Z_{s}$ and $Z_{t}$, we take the activations $Z_{s}^{(l)},Z_{t}^{(l)}$ at a certain layer $l$, project them into ${\bm{U}}_{(k)}$, and compute their inner product under the metric induced by ${\bm{S}}$:
$d(Z_{s},Z_{t})\triangleq Z^{(l)\top}_{s}{\bm{U}}_{(k)}{\bm{S}}_{(k)}^{2}{\bm{U}}_{(k)}^{\top}Z^{(l)}_{t}.$ (1)
5.3 Evaluations of the Hessian-based algorithm
Figure 4: The accuracy of swapping binding information in name (attribute) activations by projecting into ${\bm{U}}_{(k)}$ (${\bm{V}}_{(k)}$) against $k$ in a context with 3 names and 3 attributes. We test the subspaces from the Hessian (blue), a random baseline (orange), and a skyline subspace obtained by estimating the subspace spanned by the first 3 binding vectors. We perform all 3 pairwise switches: 0-1 represents swapping the binding information of $E_{0}$ and $E_{1}$ ($A_{0}$ and $A_{1}$), and so on.
While the utility of the binding matrix ${\bm{H}}$ is ultimately demonstrated in the performance of propositional probes, in this section we independently evaluate the Hessian-based algorithm. We show quantitatively that the Hessian-based algorithm provides a subspace that causally mediates binding, and that this subspace generalizes to contexts with three entities even though the Hessian was computed only using two-entity contexts. We then qualitatively evaluate our binding subspace by plotting the binding similarity (Eq. 1) for a few input contexts.
Interchange interventions   To evaluate the claim that the $k$-dimensional subspace ${\bm{U}}_{(k)}$ causally mediates binding for entities, we perform an interchange intervention (Geiger et al., 2021b) on this subspace at every layer. If ${\bm{U}}_{(k)}$ indeed carries the binding information but not any of the content information, swapping the activations in this subspace between two entities $E_{0}$ and $E_{1}$ ought to switch the bound pairs. Specifically, we perform the interventions across all layers $l=0,\dots,39$:
$Z_{E_{0}}^{(l)}$ $\leftarrow({\bm{I}}-{\bm{P}})Z_{E_{0}}^{(l)}+{\bm{P}}Z_{E_{1}}$ (2)
$Z_{E_{1}}^{(l)}$ $\leftarrow({\bm{I}}-{\bm{P}})Z_{E_{1}}^{(l)}+{\bm{P}}Z_{E_{0}}^{(l)},$ (3)
with ${\bm{P}}={\bm{U}}_{(k)}{\bm{U}}_{(k)}^{\top}$. If ${\bm{U}}_{(k)}$ correctly captures binding information, then we expect the binding information to have swapped for $E_{0}$ and $E_{1}$. We similarly test ${\bm{V}}_{(k)}$ for attributes by using $\smash{{\bm{P}}={\bm{V}}_{(k)}{\bm{V}}_{(k)}^{\top}}$ on attribute activations $\{Z_{A_{0}},Z_{A_{1}}\}$.
Baselines   We implement Distributed Alignment Search (DAS) (Geiger et al., 2024), which uses gradient descent to find a fixed dimensional subspace that enables interchange interventions between the two entities in two-entity contexts, for various choices of subspace dimension. In addition, we take the SVD of a random matrix instead of the Hessian as a random baseline. As a skyline, we evaluate the two-dimensional subspace spanned by the differences in binding vectors of the three entities. We obtain these difference vectors similarly as $\Delta_{E}$ and $\Delta_{A}$ used in the computation of the Hessian: we take samples of $Z_{E_{0}},Z_{E_{1}},$ and $Z_{E_{2}}$ across 200 contexts with different values of entities, average across the samples, and take differences between the three resultant mean activations. We consider this a skyline because this subspace is obtained from three entities’ binding vectors, whereas the Hessian and DAS are computed using contexts with only two entities.
Results   Fig. 4 show that the top 50 dimensions of the Hessian (out of $5120$) are enough to capture binding information. Moreover, despite being obtained from contexts with two entities, the subspace correctly identifies the direction of the third binding vector, in that it enables the swaps “0-2” and “1-2”. In contrast, random subspaces of all dimensions fail to switch binding information without also switching content. Further, while DAS finds subspaces that successfully swaps the first two binding vectors, they do not enable swaps between the second and third binding vectors (1-2). Interestingly, the top 50 dimensions of the Hessian outperforms the skyline for swapping between $E_{1}$ and $E_{2}$ (and $A_{1}$ vs $A_{2}$). This could be due to the fact that the skyline subspace is obtained from differences in binding vectors, which could contain systematic errors that do not contribute to binding, or perhaps due to variability in binding vectors in individual inputs. We discuss details in Appendix D.
Qualitative evaluations   We visualize the binding metric in various contexts by plotting pairwise binding similarities (Eq. 1) between token activations. Specifically, on a set of activations $Z_{0},\dots,Z_{S-1}$, we compute the matrix $M_{st}=d(Z_{s},Z_{t})$. We do so for three input contexts. We first evaluate an input context in which the entities are bound serially: the first sentence binds $E_{0}$ to $A_{0}$, and the second binds $E_{1}$ to $A_{1}$. To ensure that the binding subspace is picking up on binding, and not something spurious such as sentence boundaries, we evaluate an input context in which entities are bound in parallel: there is now only one sentence that binds $E_{0}$ and $E_{1}$ to $A_{0}$ and $A_{1}$ respectively. Lastly, we plot the similarity matrix for a context with three entities.
We find that in both the serial context (Fig. 5 left) and parallel context (Fig. 5 middle), the activations are clustered based on which entity they refer to. Interestingly, in three-entity input (Fig 5 right) the binding metric does not clearly discriminate between the second and third entities even though the interchange interventions showed that the binding subspace captures the difference in binding between them. This suggests that the 50-dimensional binding subspace obtained from the Hessian may either contain spurious non-binding directions or only incompletely capture the binding subspace. Thus, our current methods may be too noisy for contexts with more than two entities. Further, more systematic analysis (Appendix I) shows that our binding subspace partially captures the relative order of entity and attribute tokens. In Appendix E, we use similar plots to show that coreferred entities share the same binding vectors.
6 Propositional Probes Evaluations
In this section we evaluate propositional probes in standard and adversarial settings. In standard settings, propositional probes perform comparably with a prompting skyline, even on the complex, out-of-distribution para and trans datasets (Sec. 6.1). Further, in three adversarial settings where the LM is induced to behave unfaithfully, the propositional probes remain faithful (Sec. 6.2).
6.1 Standard settings
Standard setting Adversarial setting
Method Metric synth para trans synth (P) para (P) trans (P) trans (FT)
Prompting EM 1.00 (0.00) 0.93 (0.01) 0.40 (0.02) 0.07 (0.01) 0.04 (0.01) 0.06 (0.01) 0.00 (0.00)
Jaccard 1.00 (0.00) 0.98 (0.00) 0.78 (0.01) 0.49 (0.02) 0.48 (0.01) 0.51 (0.01) 0.00 (0.00)
Prop. Probes EM 0.97 (0.01) 0.55 (0.02) 0.26 (0.02) 0.98 (0.01) 0.55 (0.02) 0.24 (0.02) 0.09 (0.01)
Jaccard 0.99 (0.01) 0.90 (0.01) 0.78 (0.01) 0.99 (0.01) 0.90 (0.01) 0.76 (0.01) 0.68 (0.01)
Table 1: Exact-match accuracy (EM) and Jaccard Index of propositional probes and the prompting skyline on standard (Sec. 6.1) and adversarial settings (Sec. 6.2). Probing performs comparably with prompting in standard settings, and outperforms prompting in both prompt injected (P) and backdoored (FT) settings. Brackets show standard errors.
Prop. Probe Ablations Domain Probes
Dataset Metric Hessian DAS-50 DAS-1 random Names Food Countries Occ.
synth EM 0.97 (0.01) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.99 (0.00) 1.00 (0.00) 0.99 (0.00) 1.00 (0.00)
Jaccard 0.99 (0.01) 0.35 (0.00) 0.33 (0.00) 0.33 (0.00) - - - -
para EM 0.55 (0.02) 0.01 (0.00) 0.00 (0.00) 0.00 (0.00) 0.99 (0.00) 0.85 (0.02) 0.92 (0.01) 0.88 (0.01)
Jaccard 0.90 (0.01) 0.39 (0.01) 0.33 (0.00) 0.32 (0.00) - - - -
trans EM 0.26 (0.02) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.92 (0.01) 0.58 (0.02) 0.88 (0.01) 0.74 (0.02)
Jaccard 0.78 (0.01) 0.34 (0.01) 0.32 (0.00) 0.29 (0.00) - - - -
Table 2: Left: Exact-match accuracy (EM) and Jaccard Index of ablations to prop. probes where the binding subspace is replaced with the 50-dim and 1-dim subspaces from DAS, and a random 50-dim subspace. Right: EM of each domain probe, which is an upper bound on prop. probes.
@@ 6.1 Standard settings
Metrics   We evaluate the Jaccard index and exact-match accuracy (EM) between the ground-truth propositions and the set of propositions returned by the propositional probe. Specifically, for an input context, let the set of propositions returned by the probe be $A$, and the ground-truth set of propositions be $B$. The exact-match accuracy is the fraction of contexts for which $A=B$, and the Jaccard index is the average value of $|A\cap B|/|A\cup B|$. Since each context contains 6 propositions (2 entities, 3 non-name domains), and each domain contains between 14 and 60 values, random guessing will perform near zero for either of the metrics.
@@
Skyline and ablations   We compare propositional probes against a prompting skyline that iteratively asks the model questions about the context, while constraining answers to lie in appropriate domains. We first query the names present in the context. For each name, we query the associated value for every predicate (e.g. “What is the occupation of John?”). In addition, we ablate propositional probes by replacing the Hessian-based metric $d(\cdot,\cdot)$ with the Euclidean inner product in the 50-dim and 1-dim DAS subspaces (“DAS-50” and “DAS-1”), and a 50-dim random subspace (“random”) respectively. Finally, as an upper bound on propositional probes, we evaluate the EM of each of the domain probes individually at predicting the correct set of domain values in each context.
@@
Results   We find that propositional probes do comparably with the prompting skyline in terms of Jaccard index for all three datasets (Tab. 1), but do significantly worse for terms of EM for trans. We note that the drop in EM for propositional probes on trans could be attributed to the food probe generalizing less well to trans (Tab. 2 right). In addition, we find that ablating the Hessian-based metric $d(\cdot,\cdot)$ significantly hurts performance (Tab. 2 left), suggesting that our Hessian method plays a key role in identifying a useful binding subspace.
6.2 Adversarial settings